OpenStudy (anonymous):

simplify the radical expression. 3 sqrt 64x^12 a.4x^3 b.4x^4 c. 8x^3 d.8x^4

OpenStudy (unklerhaukus):

i'm not getting any of the options provided, ..

OpenStudy (anonymous):

ok I mustve stated it wrong, like the 3 is the small 3 up on the edge of the sqrt sign

OpenStudy (unklerhaukus):

oh ok \[\sqrt[3]{64x^{12}}=\sqrt[3]{64}\sqrt[3]{x^{12}}\\ %\qquad\qquad=3\sqrt{64}(x^{12})^{1/2}\\ %\qquad\qquad=3\sqrt{64}x^{12/2}\\ %\qquad\qquad=\\ %\qquad\qquad=\] like this?

OpenStudy (unklerhaukus):

\[\sqrt[3]{64x^{12}}=\sqrt[3]{64}\sqrt[3]{x^{12}}\\ \qquad\qquad=\sqrt[3]{64}(x^{12})^{1/3}\\ \qquad\qquad=\sqrt[3]{64}x^{12/3}\\ \qquad\qquad=\\ \qquad\qquad=\]

OpenStudy (anonymous):

yeah you got it!

OpenStudy (unklerhaukus):

do you know how to simplify the cube root of 64?

OpenStudy (anonymous):

would it be 4... cubed?

OpenStudy (unklerhaukus):

yeah 64 is the cube of 4 , so 4 is the cube root of 64

OpenStudy (unklerhaukus):

\[%\sqrt[3]{64x^{12}}=\sqrt[3]{64}\sqrt[3]{x^{12}}\\ %\qquad\qquad=\sqrt[3]{64}(x^{12})^{1/3}\\ \qquad\qquad=\sqrt[3]{4^3}x^{12/3}\\ \qquad\qquad=4x^{12/3}\\ \qquad\qquad=\]

OpenStudy (anonymous):

so it would be 4x^4

OpenStudy (unklerhaukus):

yep,

OpenStudy (anonymous):

cool, thank you!